The block of material shown below has a shear modulus of 50 GPa. Determine the shear

stress and the shear strain induced by the 2 kN shearing force.


1
Expert's answer
2021-01-20T17:45:45-0500

A block of material "a \ mm\times b\ mm\times c\ mm" has a shear modulus of 50 GPa. Determine the shear stress and the shear strain induced by the 2kN shearing force.

The shear stress is


"\tau=\dfrac{P}{A}=\dfrac{2\times10^3\ N}{(b\times10^{-3}m)(c\times10^{-3}m)}"

The shear strain is


"\gamma=\dfrac{\tau}{G}=\dfrac{2\times10^3\ N}{(b\times10^{-3}m)(c\times10^{-3}m)(50\times 10^9Pa)}"

If a block of material "150 \ mm\times 50\ mm\times 75\ mm" is given


"\tau=\dfrac{P}{A}=\dfrac{2\times10^3\ N}{(50\times10^{-3}m)(75\times10^{-3}m)}"

"=5.333\times10^5 Pa=0.533\ MPa"

"\gamma=\dfrac{\tau}{G}=\dfrac{2\times10^3\ N}{(50\times10^{-3}m)(75\times10^{-3}m)(50\times 10^9Pa)}"

"=1.067\times10^{-5}"